40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire
|
Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire
|
Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp
|
Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test
|
Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck
|
Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck
|
Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp
|
Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw
|
Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended
|
Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
|
Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended
|
Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
|
Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing
|
Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
|
Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful 
|
Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up
|
Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up 
|
Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up
|
Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
|
Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended
|
Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended
|
Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation
|
Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge
|
Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS
|
Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished
|
Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers
|
Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation
|
Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished
|
Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted
|
Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended
|
Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended
|
Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted
|
Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
|
Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO
|
Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  
|
Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic
|
Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic
|
Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic
|
Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic
|
Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics 
|
Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction
|
Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling 
|
Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed  
|
Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging
|
Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green
|
Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick 
|
Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick
|
Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned
|
Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm
|
Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 
|
Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber   
|
Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible
|
Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers
|
Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment
|
Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment 
|
Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
|
Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling
|
Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments
|
Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments
|
|
Message ID: 12381
Entry time: Fri Aug 5 18:21:28 2016
In reply to: 12377
Reply to this: 12382
|
Author: |
gautam |
Type: |
Update |
Category: |
SUS |
Subject: |
ETMX Ruby Wire Standoff GLUED |
|
|
- The latest twist in this apparently never-ending saga was that even though fine pitch balancing was achieved, the wire was out of the groove on both sides!
- I rectified this situation in the morning, did the fine pitch balancing in the afternoon
- Koji's suggestion of adjusting the OSEM holding plate totally did the trick, all four magnets are reasonably well centered relative to the vertical now...
- After the latest round of fine pitch balancing, we are now tilted in pitch backwards (i.e. towards the AR face) by <0.7mrad.
- Prior to gluing, I visually inspected the optic to check that (see attachments):
- Wires are in grooves on both sides
- Unglued ruby standoff has the correct "rotation", i.e. that the wire contacts the standoff after the groove has started, and leaves it before the groove ends, since the groove doesn't go all the way around the standoff
- Section of wire around the bottom half of the optic has no obvious kinks/other funny features
- Unglued standoff is in contact with the barrel
- All magnets are well clear of teflon in OSEM coils on both sides
- Eric also checked the frequencies of the various modes (PIT, YAW, POS and SIDE) by looking at the power spectrum of the free-swinging error signals on the coils. The pitch mode is now softer than before, at ~710mHz
- We then proceeded to glue the optic, using a needle to apply the glue (optic was clamped using face EQ stops, bottom EQ stops were not engaged as we felt this would affect the fine pitch balancing
- During the process, it looks like we may have inadvertently gotten some glue onto the wire (see attachments) - it doesn't look like any has seeped into the groove itself, but there is definitely some on the wire. We can possibly try cleaning this once the optic is out. In the worst case scenario, we will have to loop another section of wire, but the fine pitch balancing should be unaffected provided we did not perturb the optic too much
- Bob has said the large oven will be available to bake the cages on Tuesday, August 9th. By this time, we should have ETMY suspended as well (we were unable to glue the knocked off magnet on ETMY as the glass bowl we had for soaking the edge of the optic in acetone to remove the epoxy residue broke while I was assembling the various pieces of Teflon inside it. Steve is procuring a new one on Monday). It is still unclear when we can vacuum bake the two ETMs...
Attachments:
Attachment #1: Wire is in the groove in the unglued wire-standoff, groove rotation looks pretty good.
Attachment #2: Ruby standoff is sitting on the barrel of the optic (if you zoom in)
Attachment #3: Side magnet is well centered w.r.t OSEM coil
Attachment #4: UR magnet is well centered w.r.t OSEM coil
Attachment #5: UL magnet is well centered w.r.t OSEM coil
Attachment #6: LL magnet is well centered w.r.t OSEM coil
Attachment #7: LR magnet is well centered w.r.t OSEM coil
Attachment #8: Wire is in the groove in the glued Ruby standoff
Attachment #9: Standoff after gluing. 3-4 drops of epoxy are visible on the wire, but none looks to have seeped into the groove itself
Attachment #10: Side view of newly glued Ruby standoff
Attachment #11: Before and After gluing shots.
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|